# On Developers’ Self-Declaration of AI-Generated Code: An Analysis of Practices

### v1

@@ 3.2.2. Search and Filtration Process
We applied the finalized search terms to locate relevant code snippets from GitHub projects. The search results were filtered based on the programming languages used in the projects. To increase the language diversity in our dataset, we collected code files from the top 10 programming languages on GitHub  (Top, 2024): Python, Java, JavaScript, TypeScript, C, C++, Go, PHP, Ruby, and C#. Our search process consists of the following steps. First, we used the search terms we identified in Table 1 and applied GitHub’s programming language filter for the top 10 languages. In most cases (19 search terms), the search returned valid results. There were four search terms for which we were unable to retrieve any results that met our inclusion criteria (see Table 2). Specifically, we could not retrieve any results when we searched using the terms “by CodeWhisperer”, “proposed by GitHub Copilot”, “by Llama”, and “from Code Llama”. During the data labeling process, the first author performed the labeling, and then the other co-authors cross-validated the labeling results. We calculated the level of agreement between the authors using Cohen’s Kappa coefficient (Cohen, 1960). The value of Cohen’s Kappa coefficient was 0.82, indicating a high level of agreement among the authors.
@@
During this process, we applied our inclusion and exclusion criteria (see Table 2). First, we only included code snippets in our dataset when it is clearly stated in the self-declaration comments that they are AI-generated. Secondly, we only included AI-generated code snippets that belong to a repository with more than 1,000 lines of code to avoid dummy and small student practice projects (we used cloc [^1] https://github.com/AlDanial/cloc.git to calculate the lines of code in each repository).
The inclusion or exclusion criteria are provided in Table 2. Figure 3 shows an example of the inclusion criterion I1. This AI-generated code example meets our inclusion criterion 1 because the developer declared in the self-declaration comment that the GenerateMaze() function is generated by GitHub Copilot. It also meets our inclusion criterion 2 because the code is part of the repository GameAwareToys that contains 98 files with 46,198 lines of code, exceeding the 1,000 lines of code threshold. Figure 4 shows an example of the exclusion criterion E1 (see Table 2), as the self-declaration comment did not reveal if the code is AI-generated, but only contained the word “copilot”. Additionally, it belongs to the repository elysia-lambda, which has only 456 lines of code, which is less than the 1000 lines of code threshold (exclusion criterion 2).
Figure 4. Example of excluded AI-generated code demonstrating exclusion criterion E1
Next, we manually checked the source code files for self-declaration comments to determine whether they contain AI-generated code. Once the AI-generated code is identified, we manually reviewed the repositories associated with these code files. To collect a high-quality dataset, we gathered demographic information about the repositories where the AI-generated code was retrieved, including the number of commits, number of contributors, and application domain. We also collected the links to the source code files and self-declaration comments. All of this information was recorded in an MS Excel file, which is included in our dataset (Kashif et al., 2025).
@@
E1 The self-declaration comment does not clearly indicate that the code is AI-generated.
E2 The code belongs to a project likely to be a dummy or student project with less than 1,000 lines of code.
@@
We categorized the search results into two main categories: snippet-level and file-level self-declarations. Snippet-level is when the self-declaration comment is placed before the code snippet, while file-level is when the self-declaration comment is placed at the beginning or end of the source code file. We also labeled self-declaration comments in which developers’ sentiment regarding the AI-generated code is evident in the self-declaration comments. We manually categorized the comments in which developers appreciated the assistance of AI code generation tools as positive and the comments in which developers shared warnings and disclaimers about the use of AI-generated code as negative. We included all files we analyzed in our dataset (Kashif et al., 2025).
@@ 3.2.3. Data Extraction
In the data extraction, we first checked each repository description (README file) to understand the nature of the projects and the application domains. We ensured that we documented the exact line in the code file where the self-declaration comment was mentioned. We extracted the repository name, the line numbers, the number of self-declaration comments instances in a single file, and the total number of files containing AI-generated code snippets in a repository. We also extracted other demographic information, including the number of contributors, the number of commits, and the AI code generation tools used. This recorded information was constantly verified and reviewed by all co-authors. Table 3 provides an overview of our data extraction.
@@ 3.2.4. Data Analysis
We employed grounded theory to qualitatively analyze our data (Stol et al., 2016). Grounded theory is a bottom-up approach emphasizing theory generation instead of extending or verifying existing theories (Stol et al., 2016). Specifically, we employed open coding and constant comparison (Stol et al., 2016) to analyze self-declaration comments data collected from GitHub repositories.
Open Coding. We used the open coding technique by identifying and labeling essential ideas and concepts from self-declaration comments. The first author carefully read each self-declaration comment to extract codes. We iteratively refined the code as we examined more data. The other co-authors continuously reviewed the codes and shared their feedback to improve the reliability of the coding process. The analysis was done using MAXQDA [^2] https://www.maxqda.com/ to organize and manage the coding results.
As an example of open coding, we labeled the self-declaration comment, “NB: Much of this was written with Copilot! (although still had to fix a bunch of issues)” as a negative code quality indication. Similarly, we labeled the following self-declaration comment, “create a public function that handles request to the Bitcoin API, this code is generated by GitHub Copilot” as a short code explanation.
Constant Comparison. We developed high-level concepts and categories using the constant comparison method. This was performed iteratively as we compared the codes in individual self-declaration comments (comments under study) with the codes from all self-declaration comments we have encoded so far. The purpose of comparing codes against each other is to group codes with similar meanings into high-level categories. We developed a total of 7 high-level categories from the qualitative analysis of self-declaration comments. For instance, we categorized all positive and negative code quality indications mentioned in self-declaration comments into the code quality indications category. Similarly, we categorized the self-declaration comments containing various code descriptions and explanations into the code explanations category. The first author did the initial coding, while the other two co-authors provided feedback on these concepts by reviewing the coding. We resolved disagreements using the negotiated agreement approach (Campbell et al., 2013) to enhance the reliability of our data analysis.
@@ 3.3.1. Questionnaire Design
We provide an overview of our survey questionnaire in Table 4, which is also accessible online [^3] https://forms.gle/WUaFEKBPpfndLfXf8. We distributed the survey through our connections and various mailing lists from early September 2024. The survey was open for participants until mid-October 2024.
@@
Q1 Introduction Do you use AI-generated code in your projects? Descriptive Statistics
Q2 Demographic Which country are you working in? Descriptive Statistics
Q2 Demographic What is your highest educational degree? Descriptive Statistics
Q4 Demographic How many years of experience do you have? Descriptive Statistics
Q5 Demographic What is your role in your organization? Descriptive Statistics
Q6 Demographic
@@
Q7 RQ1, RQ2 Which AI-code generation tools you use in your projects? Descriptive Statistics
Q8 RQ2
@@
Q12 RQ2 What are the reasons for not self-declaring AI-generated code? Open Coding and Constant Comparison
@@
Next, we conducted a pilot run of the survey by distributing the questionnaire to a small subset of our target population. These participants were primarily practitioners with whom we had direct contact in the industry, allowing us to access them for feedback and suggestions easily. We received 11 valid responses from the participants of the pilot survey and made several revisions to the survey questionnaire to improve its structure and readability.
@@ 3.3.2. Data Collection
    Recruit practitioners from our contacts who are willing to contribute to our survey utilizing the convenience sampling method.
@@
These three strategies were designed to facilitate our data collection process while acknowledging the limitations inherent in non-probabilistic approaches. We sent invitation emails to the contributors of the repositories we collected during our mining study. We invited practitioners in our network via email and LinkedIn messages. We also invited participants through survey invitations posted on our social media profiles and various developer communities and groups on social media platforms, such as LinkedIn and Facebook.
@@ 3.3.3. Data Analysis
Validating Responses. To maintain the reliability of our data, we developed a set of inclusion and exclusion criteria to validate each survey response (as shown in Table 5).
@@
We included a prerequisite question about whether participants have used AI-generated code in practice. We received 12 responses from practitioners who declared that they do not use any AI-generated code in their projects, and thus, we excluded those responses from our analysis. We also excluded 15 other responses that were unclear or contained irrelevant answers to the survey questions.
Analyzing Responses. We analyzed all responses to the survey questionnaire in Table 4 using a combination of quantitative and qualitative analysis methods. For the first nine closed-ended questions, we employed descriptive statistics to ascertain the frequency of various responses, which provides an overview of the participants’ demographics and their usage of AI-generated code. For the remaining three open-ended questions, we utilized Grounded Theory (Stol et al., 2016), specifically open coding and constant comparative analysis. The first author carefully examined each valid response to extract meaningful insights. During this process, the author summarized the core ideas articulated in participants’ responses regarding their reasons for self-declaration of AI-generated code. For instance, in the answer to SQ11 about reasons for self-declaring AI-generated code, a participant expressed a positive view of self-declaration, stating in $R_{62}$, “It fosters trust with clients and team members by being upfront about the use of AI in the development process”. This response was coded as “fostering trust through transparency”. Another participant emphasized the importance of tracking AI-generated code, stating in $R_{69}$, “Placing comments at the exact location of AI-generated code allows for pinpointed identification of machine-generated contributions, which is essential for precise tracking and review”. This was coded as “tracking AI-generated code for review”. In contrast, some participants provided reasons for not self-declaring the use of AI in development in their answers to SQ12. A participant mentioned in $R_{46}$, “I’ll check the code over myself, the AI is just there to speed things up”, which we coded as “do not rely on AI tools but use them just for assistance”.
In this analysis, we employed constant comparisons to relate and contrast the codes identified from participants’ responses. The first author compared the codes derived from specific summarized ideas to identify patterns and similar semantic meanings. For example, responses that reflected trust, transparency, and accountability were grouped into the transparency, accountability, and ethical considerations category. Responses that emphasized the need for self-declaring AI-generated code to track and monitor it for future code reviews were categorized into the tracking and monitoring code for later reviews and debugging category. All authors were involved in the validation of the developed codes. In cases of disagreement, all authors engaged in discussions to reconcile differences and reach a consensus on the coding results.
4. Results
@@ 4.1. Mining Study Overview
Application domains. We found self-declared AI-generated code in projects belonging to various application domains, with Web applications being the most common domain with 191 (38.3%) repositories, followed by Utility tools with 162 (32.7%) repositories and Game applications with 79 (15.9%) repositories. 89 repositories were not categorized because the project descriptions were either not provided or unclear. Figure 6 shows a summary of the distribution of self-declared AI-generated code snippets across application domains.
Commit history. Figure 7 shows the commit history of the repositories that contain self-declared AI-generated code. We categorized repositories into six commit count ranges: from 1 – 50 to $\geq$ 5000. In total, we collected data from 586 repositories. 194 (33.1%) repositories are in the 100 to 500 commits range, indicating reasonable activity in these projects. Moreover, 137 repositories (23.4%) contain commit counts ranging from 1 to 50, showing certain inactivity in these projects. The projects containing more than 500 commits account for 136 (23.2%) repositories.
Figure 7. Distribution of repositories based on number of commits
@@
Figure 8. Distribution of repositories with multiple contributors
Finding 1 The collected code snippets are primarily written using ChatGPT and GitHub Copilot. Almost half of these code snippets were written in JavaScript and Python, and Web application was the dominant application domain. The repositories that contain the collected code snippets in our dataset are maintained mainly by small teams (less than five contributors, 58.0%).
@@ 4.2. Survey Demographics
Professional roles. Most participants (74, 67.2%) reported that they work in software development roles in their organizations, such as software developer, full stack developer, backend developer, and mobile application developer. Participants also work in management positions, including team leads, engineering managers, and CTOs (9 participants), as well as researchers (9 participants). Other roles represented in the survey include quality assurance and testing (5 participants), data-related positions (4 participants, e.g., data analysts and database administrators), security and cloud engineering (3 participants), and design and UI/UX (2 participants). An overview of the participants’ roles in their organizations is shown in Figure 11.
Figure 11. Overview of survey participants’ roles in their organizations
@@
Finding 2 Our survey received 111 valid responses from 26 countries. The survey participants mostly work in the industry in software development-related roles. The majority of survey participants work on web applications and prefer ChatGPT and GitHub Copilot for generating code.

### v2

@@ 3.2.2. Search and Filtration Process
We applied the finalized search terms to locate relevant code snippets from GitHub projects. The search results were filtered based on the programming languages used in the projects. To increase the language diversity in our dataset, we collected code files from the top 10 programming languages on GitHub  (Top, 2024): Python, Java, JavaScript, TypeScript, C, C++, Go, PHP, Ruby, and C#. We used the search terms we identified in Table 2 and applied GitHub’s programming language filter for the top 10 languages. In most cases (19 search terms), the search returned valid results. There were four search terms for which we were unable to retrieve any results that met our inclusion criteria (see Table 2).
We manually checked the source code files for self-declaration comments to determine whether they contain AI-generated code. During this process, we applied our inclusion and exclusion criteria (see Table 3). First, we only included code snippets in our dataset when it is clearly stated in the self-declaration comments that they are AI-generated. Secondly, we only included AI-generated code snippets that belong to a repository with more than 1,000 lines of code to avoid dummy and small student practice projects (we used cloc [^1] https://github.com/AlDanial/cloc.git to calculate the lines of code in each repository).
The inclusion or exclusion criteria are provided in Table 3. Figure 3 shows an example of AI-generated code that meets our inclusion criterion I1 because the developer declared in the self-declaration comment that the GenerateMaze() function is generated by GitHub Copilot. It also meets our inclusion criterion I2 because the code is part of the repository GameAwareToys that contains 98 files with 46,198 lines of code, exceeding the 1,000 lines of code threshold.
@@
Once the AI-generated code is identified, we manually reviewed the repositories associated with these code files. To collect a high-quality dataset, we gathered demographic information about the repositories where the AI-generated code was retrieved, including the number of commits, number of contributors, and application domain. We collected this information to illustrate the maturity and activity level of the repositories. Although this information is not directly tied to our RQs, it provides useful context for understanding the results and characteristics of our dataset, showing the relevance and representativeness of mined repositories. We also collected the links to the source code files and self-declaration comments.
@@
E1 The code belongs to a project likely to be a dummy or student project with less than 1,000 lines of code.
@@
We categorized the search results into two main categories: snippet-level and file-level self-declarations. Snippet-level is when the self-declaration comment is placed before the code snippet, while file-level is when the self-declaration comment is placed at the beginning or end of the source code file. We also labeled self-declaration comments in which developers’ sentiment regarding the AI-generated code is evident in the self-declaration comments. We manually categorized the comments in which developers appreciated the assistance of AI code generation tools as positive and the comments in which developers shared warnings and disclaimers about the use of AI-generated code as negative. All of this information was recorded in an MS Excel file, which is included in our dataset (Kashif et al., 2025).
During the data labeling process, the first author performed the labeling, and then the other co-authors cross-validated the labeling results. We calculated the agreement level between the authors using the average pairwise Cohen’s Kappa coefficient (Cohen, 1960), by computing Cohen’s Kappa for each pair of the three coders and then reporting the average value. The resulting Cohen’s Kappa coefficient was 0.82, indicating a high level of agreement among the authors.
We initially curated slightly over 1,200 code files from public GitHub repositories, and after applying our inclusion and exclusion criteria (see Table 3), we retained 613 code files from 586 unique repositories for data analysis. To ensure a diverse and non-redundant dataset for data analysis, we included a few representative files from repositories containing multiple files with AI-generated code. This was necessary as several repositories contained many files with highly similar self-declaration comments. For instance, the repository KV6002-NUMUNCHIES alone contained 65 files with similar self-declaration comments with small variations. On the contrary, for the analysis of the distribution of AI-generated code in RQ1.3, we recorded the total number of files containing self-declared AI-generated code snippets in “Total files with SDC” column (code-details.xlsx) in our dataset (Kashif et al., 2025).
@@ 3.2.3. Data Extraction
During data extraction, we first reviewed each repository’s description (README file) to understand the nature of the projects and their application domains. We also extracted other demographic information, including the number of contributors, the number of commits, and the AI code generation tools used. Secondly, we extracted the repository name, programming languages, the line numbers, the number of self-declaration comment instances in a single file, and the total number of files containing AI-generated code snippets in a repository. We ensured that we documented the exact line in the code file where the self-declaration comment was mentioned. We calculated the Lines of Code (LoC) for all 613 code files by retrieving their content using the GitHub REST API (cod, 2022) and counting the number of lines in each file. We also extracted the dates when lines containing self-declaration comments were added to code files in our dataset to analyze trends of self-declaring AI-generated code over time. This was achieved using the GitHub blame object obtained through the GitHub GraphQL API (Gra, 2025). In total, we examined 613 files and successfully collected timestamps for 601 files. The remaining 12 files were excluded because their repositories had been made private. This recorded information was constantly verified and reviewed by all co-authors. Table 4 provides an overview of our data extraction.
@@
D11 Commit date The date when self-declaration comment was added to the code. RQ1.3
D12 LoC The number of lines of code in a code file. RQ1.2
@@ 3.2.4. Data Analysis
We employed thematic analysis as our main methodology. We adopted a predominantly inductive and data-driven approach to provide a comprehensive interpretation of the data (Braun and Clarke, 2006). This method allowed us to systematically identify, analyze, and organize patterns in the self-declaration comments collected from GitHub repositories.
Initial Coding. We performed the initial coding by identifying and labeling essential ideas and concepts from self-declaration comments. The first author carefully read each self-declaration comment to extract codes. We iteratively refined the code as we examined more data. The other co-authors continuously reviewed the codes and shared their feedback to improve the reliability of the coding process. The analysis was done using MAXQDA [^2] https://www.maxqda.com/ to organize and manage the coding results.
As an example of initial coding, we labeled the self-declaration comment, “NB: Much of this was written with Copilot! (although still had to fix a bunch of issues)” as a negative code quality indication. Similarly, we labeled the following self-declaration comment, “create a public function that handles request to the Bitcoin API, this code is generated by GitHub Copilot” as a short code explanation.
Iterative Themes Development. We developed high-level themes by iteratively and constantly comparing the codes in the initial coding phase. We developed a total of 7 high-level themes from the qualitative analysis of self-declaration comments. For instance, we categorized all positive and negative code quality indications mentioned in self-declaration comments into the Code Quality Indications theme. Similarly, we categorized the self-declaration comments containing various code descriptions and explanations into the Code Explanations theme. The first author did the initial coding, while the other two co-authors provided feedback on these themes by reviewing the coding. We resolved disagreements using the negotiated agreement approach (Campbell et al., 2013) to enhance the reliability of our data analysis.
@@ 3.3.1. Questionnaire Design
We provide an overview of our survey questionnaire in Table 5, which is also accessible online [^3] https://forms.gle/WUaFEKBPpfndLfXf8. We distributed the survey through our connections and various mailing lists, specifically targeting software developers who have experience using AI code generation tools to ensure a relevant study population. The survey was open for participants until mid-October 2024.
@@
SQ1 Introduction Do you use AI-generated code in your projects? Descriptive Statistics
SQ2 Demographic Which country are you working in? Descriptive Statistics
SQ2 Demographic What is your highest educational degree? Descriptive Statistics
SQ4 Demographic How many years of experience do you have? Descriptive Statistics
SQ5 Demographic What is your role in your organization? Descriptive Statistics
SQ6 Demographic
@@
SQ7 RQ1, RQ2 Which AI-code generation tools you use in your projects? Descriptive Statistics
SQ8 RQ2
@@
SQ12 RQ2 What are the reasons for not self-declaring AI-generated code? Open Coding and Constant Comparison
@@
Next, we conducted a pilot run of the survey by distributing the questionnaire to a small subset of our target population. These participants were primarily practitioners with whom we had direct contact in the industry, allowing us to easily access them for feedback and suggestions. We received 11 valid responses from the participants of the pilot survey and made several revisions to the survey questionnaire to improve its structure and readability.
@@ 3.3.2. Data Collection
    Recruit practitioners from our contacts who are willing to contribute to our survey, utilizing the convenience sampling method.
@@
These three strategies were designed to facilitate our data collection process while acknowledging the limitations inherent in non-probabilistic approaches. We sent invitation emails to the contributors of the repositories we collected during our mining study. We invited practitioners in our network via email and LinkedIn messages. We also invited participants through survey invitations posted on our social media profiles and various developer communities and groups on social media platforms, such as LinkedIn and Facebook. Our survey received a total of 81 responses from social media (primarily LinkedIn, with some from Facebook), 36 responses from participants contacted directly, and 21 responses from contributors to the repositories of our mining study dataset. Participation was entirely voluntary, and no compensation was provided.
@@ 3.3.3. Data Analysis
Validating Responses. To maintain the reliability of our data, we developed a set of inclusion and exclusion criteria to review and validate each survey response (as shown in Table 6). Although we did not explicitly screen for AI-generated spam responses, the manual review process substantially reduced the likelihood of low-quality or automated submissions.
@@
We included a prerequisite question about whether participants have used AI-generated code in practice. We received 12 responses from practitioners who declared that they do not use any AI-generated code in their projects, and thus, we excluded those responses from our analysis. We also excluded 15 other responses that, upon analysis, were found to be unclear or contained irrelevant answers to the survey questions.
Analyzing Responses. We analyzed all responses to the survey questionnaire in Table 5 using a combination of quantitative and qualitative analysis methods. For the first nine closed-ended questions, we employed descriptive statistics to ascertain the frequency of various responses, which provides an overview of the participants’ demographics and their usage of AI-generated code. For the remaining three open-ended questions (SQ10, SQ11, SQ12), we employed thematic analysis and utilized the inductive approach for analysis (Braun and Clarke, 2006). Specifically, we first performed initial coding and then constantly compared the codes to develop high-level themes. The first author performed the initial coding by examining each valid response to extract and summarize the core ideas articulated in participants’ responses regarding their reasons to self-declare AI-generated code.
For instance, in the answer to SQ11 about reasons for self-declaring AI-generated code, a participant emphasized the importance of tracking AI-generated code, stating in $R_{69}$, “Placing comments at the exact location of AI-generated code allows for pinpointed identification of machine-generated contributions, which is essential for precise tracking and review”. This was coded as “tracking AI-generated code for review”. In contrast, some participants provided reasons for not self-declaring the use of AI in development in their answers to SQ12. A participant mentioned in $R_{46}$, “I’ll check the code over myself, the AI is just there to speed things up”, which we coded as “do not rely on AI tools but use them just for assistance”.
In this analysis, we performed iterative and constant comparisons to relate and contrast the codes identified from participants’ responses. The first author compared the codes derived from specific summarized ideas to identify patterns and similar semantic meanings. Responses that emphasized the need for self-declaring AI-generated code to track and monitor it for future code reviews were categorized into the Tracking and Monitoring Code for Later Reviews and Debugging theme. Responses where participants mentioned custom review and modification of AI-generated code as their reasons for not self-declaring it were grouped under the Customization and Modification theme. All authors were involved in the validation of the developed codes. In cases of disagreement, all authors engaged in discussions to reconcile differences and reach a consensus on the coding results.
@@ 3.4. Mining Study Overview
Application domains. We found self-declared AI-generated code in projects belonging to various application domains, with Web applications being the most common domain with 191 (32.5%) repositories, followed by Utility tools with 162 (27.7%) repositories and Game applications with 79 (13.5%) repositories. 89 repositories were not categorized because the project descriptions were either not provided or unclear. Figure 5 shows a summary of the distribution of self-declared AI-generated code snippets across application domains.
Figure 6. Distribution of repositories based on number of commits and multiple contributors
Commit history. Figure 6 shows the commit history of the repositories that contain self-declared AI-generated code. We categorized repositories into six commit count ranges: from 1 – 50 to $\geq$ 5000. In total, we collected data from 586 repositories. 194 (33.1%) repositories are in the 100 - 500 commits range, indicating reasonable activity in these projects. Moreover, 137 repositories (23.4%) contain commit counts ranging from 1 to 50, showing certain inactivity in these projects. The projects containing more than 500 commits account for 136 (23.2%) repositories.
@@
Summary of Mining Study Overview The collected code snippets are primarily written using ChatGPT and GitHub Copilot. Almost half of these code snippets were written in JavaScript and Python, and Web application was the dominant application domain. The repositories that contain the collected code snippets in our dataset are maintained mainly by small teams (less than five contributors, 58.0%).
@@ 3.5. Survey Demographics
Professional roles. Most participants (74, 67.2%) reported that they work in software development roles in their organizations. 38 participants provided a broad development role “software developer” while 36 participants mentioned specialized roles, including full-stack developer, backend developer, and mobile developer in their responses. Participants also work in management positions, including team leads, engineering managers, and CTOs (9 participants), as well as researchers (9 participants). Other roles represented in the survey include quality assurance and testing (5 participants), data-related positions (4 participants, e.g., data analysts and database administrators), security and cloud engineering (3 participants), and design and UI/UX (2 participants). An overview of the participants’ roles in their organizations is shown in Figure 9.
Figure 9. Overview of 111 total survey participants’ roles in their organizations
@@
Summary of Survey Demographics Our survey received 111 valid responses from 26 countries. The survey participants mostly work in the industry in software development-related roles. The majority of survey participants work on web applications and prefer ChatGPT and GitHub Copilot for generating code.
4. Results